# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

### v5

@@ VI Conclusion
We introduce LAMBDA (), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments. Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
To evaluate the benchmark, we tested multiple models to assess their performance and data efficiency on these complex tasks. Learning-based models, RT-1 and MotionGlot, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to ’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by ’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
Limitations and Future Work. While the inclusion of a quadruped like Spot in the real world tasks highlights the advantages of quadrupedal platforms, the lack of simulated quadruped data limits the ability to leverage diverse environments that simulators can provide, such as a broader range of floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enable greater generalization and provide richer datasets for evaluation. Furthermore, extending the benchmark to include tasks beyond pick-and-place, such as those involving more intricate manipulation, would increase its relevance and applicability to a wider range of robotic systems and models. Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency. Evaluating their performance on ’s sparse rewards would offer additional insights into their suitability for long-horizon MoMa tasks and broaden the scope of the benchmark.
| Model | Avg RoC | ||||
| RT-1 | |||||
| LSTM |

### v6

@@ VI Conclusion
We introduce LAMBDA (), a challenging benchmark consisting of a realistic amount, , of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments. Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
To evaluate the benchmark, we tested multiple baseline models to assess their performance and data efficiency on these complex tasks. End-to-end learning-based models, RT-1 and MG-MoMa, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to âs long-horizon tasks. In contrast, a neuro-symbolic system, LIMP, significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by âs tasks. We envision this benchmark as a valuable resource for the robotics community, highlighting data efficiency limitations in current models and guiding the development of more efficient systems for daily indoor tasks.
Limitations and Future Work. Incorporating a quadruped like Spot in real-world tasks showcases the advantages of quadrupedal platforms, but the absence of simulated quadruped data limits exposure to diverse environments, such as varied floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enhance generalization and enrich evaluation datasets. In the future, the benchmark could be extended to include tasks beyond pick-and-place, such as those involving more intricate manipulation.
ACKNOWLEDGMENT
This work is supported by ONR under grant award numbers N00014-22-1-2592 and N00014-23-1-2794, NSF under grant award number CNS-2150184, and with support from Amazon Robotics. We thank Aryan Singh, George Chemmala, Ziyi Yang, David Paulius, Ivy He, Lakshita Dodeja, Mingxi Jia, Benned Hedegaard, Thao Nguyen, Tuluhan Akbulut, Chaerin Min, Selena Williams, and George Konidaris.